# Diffusion approximation of the stationary distribution of a two-level single server queue

### v5

@@ 4.2 Preliminary lemmas
$g^{(r)}_{i,\theta}(x_{1},x_{2})=e^{-\eta^{(r)}_{i}}(\theta)(x_{1}\wedge 1/r)-\zeta^{(r)}(\theta)(x_{2}\wedge 1/r),\qquad(x_{1},x_{2})\in\mathbb{R}_{+}^{2},$
where $\eta^{(r)}_{i}(\theta),\zeta^{(r)}(\theta)$ are the unique solutions of the following equations for each $\theta,\theta\in\mathbb{R}$.
@@
because $\eta^{(r)}_{i}(\theta)$ is the inverse function of the Laplace transform of the finite positive random variable $T^{(r)}_{e,i}\wedge 1/r$ taking the value $e^{-\theta}$ and $\zeta^{(r)}(\theta)$ is similarly determined. Furthermore, $\eta^{(r)}_{i}(\theta),\zeta^{(r)}(\theta)$ are infinitely many differentiable functions of $\theta$. Hence, twice differentiating the equations in (4.8), the following lemma can be obtained by their Taylor expansions around the origin. It is formally proved in [3, Lemma 5.8].
@@
Furthermore, by , for $i=1,2$, there are constant $d_{e,i},d_{s},a>0$ such that, for $r\in(0,1]$ and $\theta\in\mathbb{R}$ satisfying $r|\theta|<a$,
@@
We define test function $f^{(r)}_{{\boldsymbol{\theta}}}$ for the BAR as
$f^{(r)}_{{\boldsymbol{\theta}}}(z,x_{1},x_{2})=\sum_{i=1,2}e^{\theta_{i}z}g^{(r)}_{i,\theta_{i}}(x_{1},x_{2})w^{(r)}_{i}(z),$ (4.13)
and auxiliary functions $\psi^{(r)}_{i}(\theta_{i})$ for $i=1,2$ as
@@
Based on these results, we compute $\mathbb{E}\left[{\mathcal{H}}f(X^{(r)}(0))\right]$ in (4.1) for $f^{(r)}_{r{\boldsymbol{\theta}}}$. Then,
$\mathbb{E}\left[{\mathcal{H}}f^{(r)}_{{\boldsymbol{\theta}}}(X^{(r)}(0))\right]=\sum_{i=1,2}\eta_{i}^{(r)}(r\theta_{i})\psi^{(r)}_{i}(\theta_{i})$
@@
$\mathbb{E}\left[{\mathcal{H}}f^{(r)}_{{\boldsymbol{\theta}}}(X^{(r)}(0))\right]=\frac{1}{2}\sum_{i=1,2}c_{i}\sigma_{i}^{2}\left(-\beta_{i}+\theta_{i}\right)\mu\theta_{i}r^{2}\psi^{(r)}_{i}(\theta_{i})$
$\hskip 64.58313pt+c^{(r)}_{1}\left(1-\frac{1}{2}\mu^{2}\sigma_{s}^{2}r\theta_{1}\right)\mu r\theta_{1}\mathbb{E}\left[1(L^{(r)}=0)g^{(r)}_{1,r\theta_{1}}(R^{(r)})\right]+o(r^{2}),$ (4.17)
We next compute the Palm expectation terms in (4.1). In what follows, we use the following simplified notations.
@@
Since $g^{(r)}_{i,\theta_{i}}$ changes only when $L^{(r)}(0-)=\ell^{(r)}_{0}$ under $\mathbb{P}^{(r)}_{e}$ and $L^{(r)}(0)=\ell^{(r)}_{0}$ under $\mathbb{P}^{(r)}_{s}$, (4.8) implies that
$\mathbb{E}^{(r)}_{e}\left[\Delta_{e}f^{(r)}_{{\boldsymbol{\theta}},r}(X^{(r)}(0))1(L^{(r)}(0-)\not=\ell^{(r)}_{0})\right]=0,$
$\mathbb{E}^{(r)}_{s}\left[\Delta_{s}f^{(r)}_{{\boldsymbol{\theta}},r}(X^{(r)}(0))1(L^{(r)}(0)\not=\ell^{(r)}_{0})\right]=0.$
@@
$(k+1)(\min_{i=1,2}c^{(r)}_{i})\mathbb{E}[(R^{(r)}_{s})^{k}1(R^{(r)}_{s})\in B,L^{(r)}\geq 1)]\leq\alpha^{(r)}_{e}\mathbb{E}[(T^{(r)}_{s})^{k+1}1(T^{(r)}_{s}\in B)],$ (4.23)
@@
$\lim_{r\downarrow 0}\Big{(}c_{1}\mu\mathbb{E}[1(L^{(r)}=0)g^{(r)}_{1,r\theta}(R^{(r)})]+\alpha^{(r)}_{e}e^{r\theta\ell_{0}}(\lambda_{2}-\lambda_{1})\mathbb{E}_{s}[\widehat{R}^{(r)}_{e}1(L^{(r)}(0)=\ell^{(r)}_{0})]\Big{)}=0.$
Since $e^{r\theta\ell_{0}}$ can be arbitrarily small, this implies that
@@
$\mathbb{E}^{(r)}_{e}[\widehat{R}^{(r)}_{s}1(L^{(r)}(0-)=\ell^{(r)}_{0})]=O(r),\qquad\mathbb{E}^{(r)}_{s}[\widehat{R}^{(r)}_{e}1(L^{(r)}(0)=\ell^{(r)}_{0})]=O(r),$ (4.44)
@@ 5.1 Process limit
$\Big{(}\!-b_{1}\theta_{1}+\frac{1}{2}c_{1}\sigma_{1}^{2}\theta_{1}^{2}\Big{)}\widetilde{\varphi}_{1}(\theta_{1})+\Big{(}\!-b_{2}\theta_{2}+\frac{1}{2}c_{2}\sigma_{2}^{2}\theta_{2}^{2}\Big{)}\widetilde{\varphi}_{2}(\theta_{2})$
$\;\;+\left(\frac{\beta_{1}e^{\beta_{1}\ell_{0}}}{e^{\beta_{1}\ell_{0}}-1}-\frac{\beta_{1}e^{\theta_{1}\ell_{0}}}{e^{\beta_{1}\ell_{0}}-1}\right)\theta_{1}\frac{c_{1}\sigma^{2}_{1}}{2}\widetilde{\varphi}_{1}(0)+\beta_{2}e^{\theta_{2}\ell_{0}}\theta\frac{c_{2}\sigma^{2}_{2}}{2}\widetilde{\varphi}_{2}(0)=0,\quad\theta\in\mathbb{R}\times\mathbb{R}_{-},$ (5.1)
@@
We like to identify this stationary distribution $\widetilde{\nu}$ by that of a reflecting diffusion process. Assume that the diffusion scaled process $L^{(r)}_{\rm dif}(\cdot)\equiv\{rL^{(r)}(r^{-2}t);t\geq 0\}$ weakly converges to a nonnegative valued stochastic process $Z(\cdot)\equiv\{Z(t);t\geq 0\}$ as $r\downarrow 0$. We aim to characterize this $Z(\cdot)$ by a stochastic integral equation. To this end, define real-valued functions $\sigma(x)$ and $b(x)$ for $x\in\mathbb{R}_{+}$ as
$\sigma(x)=\sqrt{c_{1}}\sigma_{1}1(x\leq\ell_{0})+\sqrt{c_{2}}\sigma_{2}1(x>\ell_{0}),$
$b(x)=b_{1}1(x\leq\ell_{0})+b_{2}1(x>\ell_{0}),$
@@
So far, in the original submission of this paper, we assumed the existence of the solution $Z(\cdot)$ of (5.2), and considered the stationary distribution using the generalized Ito formula for a convex function. We got a stationary equation similar to (5.1) for the $Z(\cdot)$, but could not completely obtained the stationary distribution although we had partial information which is compatible with Theorem 3.1. So, it was conjectured that
@@
During the reviewing process of this paper, the author have known that (a) is resolved by Atar et al. [1, Lemma 4.3] (see also [12] for a different proof), while the author affirmatively answers (c) in [12]. Those results are obtained for more general reflecting diffusion processes than $Z(\cdot)$, which are hard to include in this paper. Thus, we refer to [12] for them, and will not further discuss about the conjectures (a) and (c). As for the conjecture (b), it is still open (see [12]).
@@
We intuitively show (3.12). First assume that $b_{1}\not=0$. Under the stationary setting, let $\tau^{(r)}_{1}$ be the first time for $L^{(r)}(t)$ to get into $\ell^{(r)}_{0}$ from below or above. Suppose that $L^{(r)}(t)$ first moves to $\ell^{(r)}_{0}+1$ after time $\tau^{(r)}_{1}$, and returns to $\ell^{(r)}_{0}$ after time $B^{(r)}_{2}$. Since the relative frequency of moving up at level $\ell^{(r)}_{0}$ is guessed to be $\lambda^{(r)}_{2}/(\lambda^{(r)}_{2}+c^{(r)}_{1}\mu^{(r)})$ and $B^{(r)}_{2}$ is stochastically almost identical with the busy period of the $GI/G/1$ queue with inter-arrival time $T^{(r)}_{e,2}$ and service time $T^{(r)}_{s}/c^{(r)}_{2}$, we have
$\mathbb{E}[B^{(r)}_{2}]\sim\frac{\lambda^{(r)}_{2}}{\lambda^{(r)}_{2}+c^{(r)}_{1}\mu^{(r)}}\times\frac{1}{c^{(r)}_{2}\mu^{(r)}(1-\rho^{(r)}_{2})}=\frac{c_{2}}{c_{1}+c_{2}}\times\frac{1}{rb_{2}\mu}+o(1),$ (5.3)
@@
On the other hand, when $L^{(r)}(t)$ first moves to $\ell^{(r)}_{0}-1$ after time $\tau^{(r)}_{1}$, let $B^{(r)}_{1}$ be the time to returns to $\ell^{(r)}_{0}$ from below. Then, similarly to $B^{(r)}_{2}$ but, taking $b_{1}\not=0$ and the reflecting boundary at level 00 into account, it follows from (4.25) of [11] that
$\mathbb{E}[B^{(r)}_{1}]\sim\frac{c^{(r)}_{1}\mu^{(r)}}{\lambda^{(r)}_{2}+c^{(r)}_{1}\mu^{(r)}}\times\frac{1}{\lambda^{(r)}_{1}\mathbb{P}^{(r)}_{s}[L^{(r)}(0)=\ell^{(r)}_{0}]}=\frac{c_{1}}{c_{1}+c_{2}}\times\frac{e^{\beta_{1}\ell_{0}}-1}{rb_{1}\mu}+o(1).$ (5.4)
Since $B^{(r)}_{2}+B^{(r)}_{1}$ is the cycle time to return to $\ell^{(r)}_{0}$, it follows from the cycle formula of a regenerative process that
$\lim_{r\downarrow 0}\nu^{(r)}_{1}(\mathbb{R}_{+})=\lim_{r\downarrow 0}\frac{\mathbb{E}[B^{(r)}_{1}]}{\mathbb{E}[B^{(r)}_{1}]+\mathbb{E}[B^{(r)}_{2}]}=\frac{c_{1}b_{2}(e^{\beta_{1}\ell_{0}}-1)}{c_{2}b_{1}+c_{1}b_{2}(e^{\beta_{1}\ell_{0}}-1)}=\widetilde{A}_{1}.$
Thus, for $b_{1}\not=0$, (3.12) is intuitively obtained for $i=1$, and it is similarly obtained for $i=2$. The case for $b_{1}=0$ is similarly derived. Thus, we have another support of (3.12).
@@
We prove Proposition 3.1 only for $b_{1}\not=0$ because the case for $b_{1}=0$ can be similarly proved. We first consider this 2-level $M/M/1$ queue without index $r$ for notational convenience. Because of the exponential assumptions, we can describe this model by the continuous time Markov chain with state space
$S=\{(m,i);m=1,2,i=0,1,\ldots,\ell_{0}\}\cup\{(3,i);i\geq\ell_{0}+1\},$
and its transition diagram is depicted in Figure A.1.
@@
Denote this Markov chain by $X_{\exp}(\cdot)\equiv\{(J(t),L(t));t\geq 0\}$. Note that arrivals may find the queue length less than $\ell_{0}$ when their last arrivals found the queue length not less than $\ell_{0}$ (see the arrival at time $t^{(r)}_{e}(8)$ in Figure 3.1). Because of this, the arrival rate at any time may depend on the queue length at the last arrival instant, and therefore $L(\cdot)$ cannot be a continuous-time Markov chain.
We partition the state space $S$ of $X_{\exp}(\cdot)$ to the following three disjoint subsets.
$S_{m}=\{(m,i)\in S;i\in[0,\ell_{0}]\},\quad m=1,2,\qquad S_{3}=\{(3,i):i\geq\ell_{0}+1\}.$
By the assumptions of Proposition 3.1, $X_{M}(\cdot)$ has a stationary distribution. Let $(J,L)$ be a random vector subject to this stationary distribution, and define $p_{m,i}$ as
$p_{m,i}=\mathbb{P}[J=m,L=i],\qquad m=1,2,3,(m,i)\in S_{m}.$
Obviously, the stationary distribution, $\{p_{m,i};(m.i)\in S\}$, is uniquely determined by the following stationary equations. For state in $S_{1}$,
@@
$(\lambda_{1}+c_{1}\mu)p_{1,i}=\lambda_{1}p_{1,i-1}+c_{1}\mu p_{1,i+1}+\lambda_{2}p_{2,i-1},\quad i\in[1,\ell_{0}-1],$ (A.2)
@@
For state in $S_{3}$, it follows from the equation for $\ell_{0}+1$ and the local balance that
@@
$\rho_{ij}=\frac{\lambda_{i}}{c_{j}\mu},\qquad\gamma_{ij}=\frac{c_{i}\mu}{c_{1}\mu+\lambda_{j}},\qquad i,j=1,2.$
@@
$p_{3,i}=\rho_{22}^{i-(\ell_{0}+1)}p_{3,\ell_{0}+1},\qquad i\geq\ell_{0}+1.$ (A.9)
Thus, $p_{3,i}$ for $i\geq\ell_{0}+1$ is simply determined by $p_{3,\ell_{0}+1}$.
We next consider $p_{2,i}$ for $i\in[0,\ell_{0}-1]$. From (A.4) and (A.5), we have
@@
$p_{2,i}=\gamma_{12}p_{2,i+1}=\gamma_{12}^{\ell_{0}-i}p_{2,\ell_{0}},\quad i\in[1,\ell_{0}-1].$ (A.11)
Thus, $p_{2,i}$ for $i\in[0,\ell_{0}-1]$ is determined by $p_{2,\ell_{0}}$. Furthermore, (A.11) can be written as
$p_{2,i}=\gamma_{12}^{-1}p_{2,i-1}=\gamma_{12}^{-i+1}p_{2,1}=\gamma_{12}^{-i+1}\rho_{21}p_{2,0},\quad i\in[1,\ell_{0}-1].$ (A.12)
Let us consider relations among the stationary probabilities, $p_{1,0},p_{1,\ell_{0}}$, $p_{2,\ell_{0}}$ and $p_{3,\ell_{0}+1}$ at the boundaries states of $S_{m}$ for $m=1,2,3$. We call them boundary probabilities. There is one more boundary probability, $p_{2,0}$, but this probability is ignorable compared with $p_{2,\ell_{0}}$ for large $\ell_{0}$ as $\rho_{21}p_{2,0}=\gamma_{12}^{\ell_{0}-1}p_{2,\ell_{0}-1}=\gamma_{12}^{\ell_{0}}p_{2,\ell_{0}}$ by (A.12) and (A.11), so we do not consider it. We first note that, from (A.6),
@@
Since $c_{2}\mu p_{\ell_{0}+2}=\lambda_{2}p_{3,\ell_{0}+1}$ by (A.8) for $i=\ell_{0}+1$, (A.7) can be written as
@@
To resolve this problem, we consider $p_{1,i}$ for $i\in[0,\ell_{0}-1]$. From (A.2),
$\lambda_{1}p_{1,i}-c_{1}\mu p_{1,i+1}=\lambda_{1}p_{1,i-1}-c_{1}\mu p_{1,i}+\lambda_{2}p_{2,i-1},\qquad i\in[1,\ell_{0}-1],$

### v6

@@ 4.2 Preliminary lemmas
$g^{(r)}_{i,\theta}(x_{1},x_{2})=\exp\left(-\eta^{(r)}_{i}(\theta)(x_{1}\wedge 1/r)-\zeta^{(r)}(\theta)(x_{2}\wedge 1/r)\right),\qquad(x_{1},x_{2})\in\mathbb{R}_{+}^{2},$
where $\eta^{(r)}_{i}(\theta),\zeta^{(r)}(\theta)$ are the unique solutions of the following equations for each $\theta\in\mathbb{R}$.
@@
because $\eta^{(r)}_{i}(\theta)$ is the inverse function of the Laplace transform of the finite positive random variable $T^{(r)}_{e,i}\wedge 1/r$ taking the value $e^{-\theta}$ and $\zeta^{(r)}(\theta)$ is similarly determined. Then, for each ${\boldsymbol{\theta}}\equiv(\theta_{1},\theta_{2})\in\mathbb{R}^{2}$, define test function $f^{(r)}_{{\boldsymbol{\theta}}}$ for the BAR as
$f^{(r)}_{{\boldsymbol{\theta}}}(z,x_{1},x_{2})=\sum_{i=1,2}e^{\theta_{i}z}g^{(r)}_{i,\theta_{i}}(x_{1},x_{2})w^{(r)}_{i}(z),\qquad(z,x_{1},x_{2})\in S.$ (4.9)
Functions $\eta^{(r)}_{i}(\theta)$ and $\zeta^{(r)}(\theta)$ play key roles in our analysis. Consider their properties. First, it is easy to see that they are infinitely many differentiable functions of $\theta$. Hence, twice differentiating the equations in (4.8), the following lemma can be obtained by their Taylor expansions around the origin. It is formally proved in [3, Lemma 5.8].
@@
Furthermore, for $i=1,2$, there are constant $d_{e,i},d_{s},a>0$ such that, for $r\in(0,1]$ and $\theta\in\mathbb{R}$ satisfying $r|\theta|<a$,
@@
We next introduce auxiliary functions $\psi^{(r)}_{i}(\theta_{i})$ for $i=1,2$ as
@@
Based on these facts, we compute $\mathbb{E}\left[{\mathcal{H}}f(X^{(r)}(0))\right]$ in (4.1) for $f^{(r)}_{r{\boldsymbol{\theta}}}$. Then,
$\mathbb{E}\left[{\mathcal{H}}f^{(r)}_{r{\boldsymbol{\theta}}}(X^{(r)}(0))\right]=\sum_{i=1,2}\eta_{i}^{(r)}(r\theta_{i})\psi^{(r)}_{i}(\theta_{i})$
@@
$\mathbb{E}\left[{\mathcal{H}}f^{(r)}_{r{\boldsymbol{\theta}}}(X^{(r)}(0))\right]=\frac{1}{2}\sum_{i=1,2}c_{i}\sigma_{i}^{2}\left(-\beta_{i}+\theta_{i}\right)\mu\theta_{i}r^{2}\psi^{(r)}_{i}(\theta_{i})$
$\qquad+c^{(r)}_{1}\left(1-\frac{1}{2}\mu^{2}\sigma_{s}^{2}r\theta_{1}\right)\mu r\theta_{1}\mathbb{E}\left[1(L^{(r)}=0)g^{(r)}_{1,r\theta_{1}}(R^{(r)})\right]+o(r^{2}).$ (4.17)
We next compute the Palm expectation terms in (4.1). In what follows, we use the following abbreviated notations.
@@
Since $g^{(r)}_{i,\theta_{i}}$ changes only when $L^{(r)}(0-)=\ell^{(r)}_{0}$ under $\mathbb{P}^{(r)}_{e}$ or $L^{(r)}(0)=\ell^{(r)}_{0}$ under $\mathbb{P}^{(r)}_{s}$, (4.8) implies that
$\mathbb{E}^{(r)}_{e}\left[\Delta_{e}f^{(r)}_{r{\boldsymbol{\theta}}}(X^{(r)}(0))1(L^{(r)}(0-)\not=\ell^{(r)}_{0})\right]=0,$
$\mathbb{E}^{(r)}_{s}\left[\Delta_{s}f^{(r)}_{r{\boldsymbol{\theta}}}(X^{(r)}(0))1(L^{(r)}(0)\not=\ell^{(r)}_{0})\right]=0.$
@@
$(k+1)(\min_{i=1,2}c^{(r)}_{i})\mathbb{E}[(R^{(r)}_{s})^{k}1(R^{(r)}_{s})\in B,L^{(r)}\geq 1)]$
$\qquad\leq\alpha^{(r)}_{e}\mathbb{E}[(T^{(r)}_{s})^{k+1}1(T^{(r)}_{s}\in B)],$ (4.23)
@@
$\lim_{r\downarrow 0}\Big{(}c_{1}\mu\mathbb{E}[1(L^{(r)}=0)g^{(r)}_{1,r\theta}(R^{(r)})]+\alpha^{(r)}_{e}e^{r\theta\ell^{(r)}_{0}}(\lambda_{2}-\lambda_{1})\mathbb{E}_{s}[\widehat{R}^{(r)}_{e}1(L^{(r)}(0)=\ell^{(r)}_{0})]\Big{)}=0.$
Since $e^{r\theta\ell^{(r)}_{0}}=e^{(\ell_{0}+o(1))\theta}$ can be arbitrarily small by choosing small $\theta<0$, we have
@@
$\mathbb{E}^{(r)}_{e}[\widehat{R}^{(r)}_{s}1(L^{(r)}(0-)=\ell^{(r)}_{0})]=O(r),\quad\mathbb{E}^{(r)}_{s}[\widehat{R}^{(r)}_{e}1(L^{(r)}(0)=\ell^{(r)}_{0})]=O(r),$ (4.44)
@@ 5.1 Process limit
$\sum_{i=1}^{2}\Big{(}\!-b_{i}\theta_{i}+\frac{1}{2}c_{i}\sigma_{i}^{2}\theta_{i}^{2}\Big{)}\widetilde{\varphi}_{i}(\theta_{i})+\left(\frac{\beta_{1}e^{\beta_{1}\ell_{0}}}{e^{\beta_{1}\ell_{0}}-1}-\frac{\beta_{1}e^{\theta_{1}\ell_{0}}}{e^{\beta_{1}\ell_{0}}-1}\right)\theta_{1}\frac{c_{1}\sigma^{2}_{1}}{2}\widetilde{\varphi}_{1}(0)$
$\qquad+\beta_{2}e^{\theta_{2}\ell_{0}}\theta_{2}\frac{c_{2}\sigma^{2}_{2}}{2}\widetilde{\varphi}_{2}(0)=0,\qquad\theta\equiv(\theta_{1},\theta_{2})\in\mathbb{R}\times\mathbb{R}_{-},$ (5.1)
@@
We like to identify this stationary distribution $\widetilde{\nu}$ by that of a reflected diffusion process. Assume that the diffusion scaled process $L^{(r)}_{\rm dif}(\cdot)\equiv\{rL^{(r)}(r^{-2}t);t\geq 0\}$ weakly converges to a nonnegative valued stochastic process $Z(\cdot)\equiv\{Z(t);t\geq 0\}$ as $r\downarrow 0$. We aim to characterize this $Z(\cdot)$ by a stochastic integral equation. To this end, define real-valued functions $\sigma(x)$ and $b(x)$ for $x\in\mathbb{R}_{+}$ as
$b(x)=b_{1}\mu 1(x\leq\ell_{0})+b_{2}\mu 1(x>\ell_{0}),$
$\sigma(x)=\sqrt{c_{1}}\mu\sigma_{1}1(x\leq\ell_{0})+\sqrt{c_{2}}\mu\sigma_{2}1(x>\ell_{0}),$
@@
In the original submission of this paper, we assumed the existence of the solution $Z(\cdot)$ of (5.2), and considered its stationary distribution using the Ito formula, which yields
$\sum_{i=1}^{2}\Big{(}\!-b_{i}\theta+\frac{1}{2}c_{i}\sigma_{i}^{2}\theta^{2}\Big{)}\widehat{\varphi}_{i}(\theta)+\frac{\beta_{1}e^{\beta_{1}\ell_{0}}}{e^{\beta_{1}\ell_{0}}-1}\theta\frac{c_{1}\sigma^{2}_{1}}{2}\widehat{\varphi}_{1}(0)=0,\qquad\theta\in\mathbb{R}.$ (5.3)
where $\widehat{\varphi}_{1}(\theta)=\mathbb{E}[e^{\theta Z}1(Z\leq\ell_{0})]$ and $\widehat{\varphi}_{2}(\theta)=\mathbb{E}[e^{\theta Z}1(Z>\ell_{0})]$ for random variable $Z$ subject to the stationary distribution of the $Z(\cdot)$. However, we had not solved this stationary equation. So, it was conjectured that
@@
During the review process of this paper, the author have known that (a) is resolved by Atar et al. [1, Lemma 4.3] (see also Miyazawa [11] for a different proof), while (c) can be well answered because the stationary distribution of the $Z(\cdot)$ is fully obtained in [11], in which it is claimed that the limiting distribution $\widetilde{\nu}$ of this paper is identical with the stationary distribution of the $Z(\cdot)$, but this claim is only true if and only if $c_{1}=c_{2}$. This fact also can be seen from the comparison of (5.3) with (5.1), which does not require the $\widetilde{\nu}$ itself. However, the $\widetilde{\nu}$ has not yet been completely specified.
As for (a), the stationary distributions are obtained for more general reflected diffusion processes than $Z(\cdot)$. We refer to [11, 13] for them, and will not further discuss about the conjectures (a) and (c). As for the conjecture (b), it is still open.
@@
We intuitively show (3.12). First assume that $b_{1}\not=0$. Under the stationary setting, let $\tau^{(r)}_{1}$ be the first time for $L^{(r)}(t)$ to get into $\ell^{(r)}_{0}$ from below or above. Suppose that $L^{(r)}(t)$ first moves to $\ell^{(r)}_{0}+1$ after time $\tau^{(r)}_{1}$, and returns to $\ell^{(r)}_{0}$ at time $\tau^{(r)}_{2}$ after time $B^{(r)}_{2}$. Since the relative frequency of moving up at level $\ell^{(r)}_{0}$ is guessed to be $f^{(r)}_{\rm up}\equiv\lambda^{(r)}_{2}/(\lambda^{(r)}_{2}+c^{(r)}_{1}\mu^{(r)})$ and $B^{(r)}_{2}$ is stochastically almost identical with the busy period of the $GI/G/1$ queue with inter-arrival time $T^{(r)}_{e,2}$ and service time $T^{(r)}_{s}/c^{(r)}_{2}$, the mean sojourn time for $L^{(r)}(t)$ to be above $\ell^{(r)}_{0}$ in the cycle $(\tau^{(r)}_{1},\tau^{(r)}_{2}]$ is
$f^{(r)}_{\rm up}\mathbb{E}[B^{(r)}_{2}]\sim\frac{\lambda^{(r)}_{2}}{\lambda^{(r)}_{2}+c^{(r)}_{1}\mu^{(r)}}\times\frac{1}{c^{(r)}_{2}\mu^{(r)}(1-\rho^{(r)}_{2})}=\frac{c_{2}}{c_{1}+c_{2}}\times\frac{1}{rb_{2}\mu}+o(1),$ (5.4)
@@
On the other hand, when $L^{(r)}(t)$ first moves to $\ell^{(r)}_{0}-1$ after time $\tau^{(r)}_{1}$, let $B^{(r)}_{1}$ be the time to returns to $\ell^{(r)}_{0}$ from below. Let $f^{(r)}_{\rm down}=1-f^{(r)}_{\rm up}$, then, similarly to $B^{(r)}_{2}$ but, taking $b_{1}\not=0$ and the reflecting boundary at level 00 into account, it follows from (4.25) of [12] that the mean sojourn time for $L^{(r)}(t)$ to be below $\ell^{(r)}_{0}$ in the cycle $(\tau^{(r)}_{1},\tau^{(r)}_{2}]$ is
$f^{(r)}_{\rm down}\mathbb{E}[B^{(r)}_{1}]$ $\sim\frac{c^{(r)}_{1}\mu^{(r)}}{\lambda^{(r)}_{2}+c^{(r)}_{1}\mu^{(r)}}\times\frac{1}{\lambda^{(r)}_{1}\mathbb{P}^{(r)}_{s}[L^{(r)}(0)=\ell^{(r)}_{0}]}$
$=\frac{c_{1}}{c_{1}+c_{2}}\times\frac{e^{\beta_{1}\ell_{0}}-1}{rb_{1}\mu}+o(1).$ (5.5)
Since $f^{(r)}_{\rm up}B^{(r)}_{2}+f^{(r)}_{\rm down}B^{(r)}_{1}$ is the mean cycle time to return to $\ell^{(r)}_{0}$, it follows from the cycle formula of a regenerative process that
$\lim_{r\downarrow 0}\nu^{(r)}_{1}(\mathbb{R}_{+})=\lim_{r\downarrow 0}\frac{f^{(r)}_{\rm down}\mathbb{E}[B^{(r)}_{1}]}{f^{(r)}_{\rm down}\mathbb{E}[B^{(r)}_{1}]+f^{(r)}_{\rm up}\mathbb{E}[B^{(r)}_{2}]}=\frac{c_{1}b_{2}(e^{\beta_{1}\ell_{0}}-1)}{c_{2}b_{1}+c_{1}b_{2}(e^{\beta_{1}\ell_{0}}-1)}=\widetilde{A}_{1}.$
Thus, for $b_{1}\not=0$, (3.12) is intuitively obtained for $i=1$, and it is similarly obtained for $i=2$. The case for $b_{1}=0$ is similarly derived. Thus, we have support of (3.12).
@@
We prove Proposition 3.1 only for $b_{1}\not=0$ because the case for $b_{1}=0$ can be similarly proved. We first consider this 2-level $M/M/1$ queue without index $r$ for notational convenience. Because of the exponential assumption, this model can be described by a continuous-time Markov chain. However, $L(\cdot)$ itself cannot be a Markov chain because the arrival rate may depend on the queue length at the last arrival instant as discussed at the end of Section 3.2. Taking this into account, we describe this model by the continuous time Markov chain with state space
$S=\{(j,\ell);j=1,2,\ell=0,1,\ldots,\ell_{0}\}\cup\{(3,\ell);\ell\geq\ell_{0}+1\},$
and its transition diagram is depicted in Figure A.1. Denote this Markov chain by $X_{\exp}(\cdot)\equiv\{(J(t),L(t));t\geq 0\}$.
@@
We partition the state space $S$ of $X_{\exp}(\cdot)$ into the following three disjoint subsets.
$S_{j}=\{(j,\ell)\in S;\ell\in[0,\ell_{0}]\},\quad j=1,2,\qquad S_{3}=\{(3,\ell):\ell\geq\ell_{0}+1\}.$
By the assumptions of Proposition 3.1, $X_{M}(\cdot)$ has a stationary distribution. Let $(J,L)$ be a random vector subject to this stationary distribution, and define $p_{j,\ell}$ as
$p_{j,\ell}=\mathbb{P}[J=j,L=\ell],\qquad j=1,2,3,(j,\ell)\in S_{j}.$
Obviously, the stationary distribution, $\{p_{j,\ell};(j,\ell)\in S\}$, is uniquely determined by the following stationary equations. For state in $S_{1}$,
@@
$(\lambda_{1}+c_{1}\mu)p_{1,\ell}=\lambda_{1}p_{1,\ell-1}+c_{1}\mu p_{1,\ell+1}+\lambda_{2}p_{2,\ell-1},\quad\ell\in[1,\ell_{0}-1],$ (A.2)
@@
For states in $S_{3}$, it follows from the equation for $(3,\ell_{0}+1)$ and the local balance that
@@
$\rho_{j,k}=\frac{\lambda_{j}}{c_{k}\mu},\qquad\gamma_{j,k}=\frac{c_{j}\mu}{c_{1}\mu+\lambda_{k}},\qquad j,k=1,2.$
@@
$p_{3,\ell}=\rho_{22}^{\ell-(\ell_{0}+1)}p_{3,\ell_{0}+1},\qquad\ell\geq\ell_{0}+1.$ (A.9)
Thus, $p_{3,\ell}$ for $\ell\geq\ell_{0}+1$ is simply determined by $p_{3,\ell_{0}+1}$.
We next consider $p_{2,\ell}$ for $\ell\in[0,\ell_{0}-1]$. From (A.4) and (A.5), we have
@@
$p_{2,\ell}=\gamma_{12}p_{2,\ell+1}=\gamma_{12}^{\ell_{0}-\ell}p_{2,\ell_{0}},\quad\ell\in[1,\ell_{0}-1].$ (A.11)
Thus, $p_{2,\ell}$ for $\ell\in[0,\ell_{0}-1]$ is determined by $p_{2,\ell_{0}}$. Furthermore, (A.11) can be written by (A.4) as
$p_{2,\ell}=\gamma_{12}^{-1}p_{2,\ell-1}=\gamma_{12}^{-\ell+1}p_{2,1}=\gamma_{12}^{-\ell+1}\rho_{21}p_{2,0},\quad\ell\in[1,\ell_{0}-1].$ (A.12)
Let us consider relations among the stationary probabilities, $p_{1,0},p_{1,\ell_{0}}$, $p_{2,\ell_{0}}$ and $p_{3,\ell_{0}+1}$ at the boundaries states of $S_{j}$ for $j=1,2,3$. We call them boundary probabilities. There is one more boundary probability, $p_{2,0}$, but this probability is ignorable compared with $p_{2,\ell_{0}}$ for large $\ell_{0}$ as $\rho_{21}p_{2,0}=\gamma_{12}^{\ell_{0}-1}p_{2,\ell_{0}-1}=\gamma_{12}^{\ell_{0}}p_{2,\ell_{0}}$ by (A.12) and (A.11), so we do not consider it. We first note that, from (A.6),
@@
Since $c_{2}\mu p_{\ell_{0}+2}=\lambda_{2}p_{3,\ell_{0}+1}$ by (A.8) for $\ell=\ell_{0}+1$, (A.7) can be written as
@@
To resolve this problem, we consider $p_{1,\ell}$ for $\ell\in[0,\ell_{0}-1]$. From (A.2),
$\lambda_{1}p_{1,\ell}-c_{1}\mu p_{1,\ell+1}=\lambda_{1}p_{1,\ell-1}-c_{1}\mu p_{1,\ell}+\lambda_{2}p_{2,\ell-1},\qquad\ell\in[1,\ell_{0}-1],$